OpenStudy (anonymous):

A sailmaker is making a right triangular sail. One of the angles in the sail is 47°. What is the measure of the unknown angle?

OpenStudy (disco619):

Dude this is so easy! Wait a sec lemme draw!

OpenStudy (disco619):

The draw button is not working so.... The sum of all angles in a triangle is = 180 Right angle triangle has one 90 angle. The second angle is given = 47° So we will take unknown angle = x 90 + 47 + x = 180 137 + x = 180 x = 180 - 137 x = 43°